OpenStudy (anonymous):

Solve the system of equations. 2x-3y+z=-1 3x+2y+2z=-1 x-y-3z=-4

OpenStudy (anonymous):

this stuff is a real pain let the machine do it

OpenStudy (anonymous):

you know how?

OpenStudy (anonymous):

Not a clue

OpenStudy (anonymous):

Which machine?

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

the machine you are typing on

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

I have another question

OpenStudy (anonymous):

The graph of which quadratic equation has a vertex of (-2,1). Y=(x+1)^2-2 Y=(x-1)^2-2 Y=(x+2)^2+1 Y=(x-2)^2+1

OpenStudy (solomonzelman):

need it by hand (for question 1), then I would suggest matrix. it is already in order for a matrix.

OpenStudy (anonymous):

I got the answer to the first question. And I don't know matrix

OpenStudy (solomonzelman):

oh.

OpenStudy (solomonzelman):

`y=(x-h)²+k` has a vertex (h,k)

OpenStudy (solomonzelman):

now, where would this (h,k) be (-2,1) ? In which of the options?

OpenStudy (anonymous):

Y=(x-1)^2+1

OpenStudy (solomonzelman):

y=(x-h)²+k you need vertex (-2,1) and in that form the vertex is (h,k) so h=-2, k=1 y=(x- (-2) )² + (1) y= ? (you tell me....)

OpenStudy (anonymous):

Um I'm not really sure

OpenStudy (misty1212):

HI!!

OpenStudy (solomonzelman):

Hello :)

OpenStudy (misty1212):

where did your \(-1\) come from? \[y=(x-\color{blue}h)^2+\color{red}k\] put \[\color{blue}h=\color{blue}{-2}\] and \[\color{red}k=\color{red}1\]

OpenStudy (anonymous):

My bad I meant to put 2 not 1

OpenStudy (misty1212):

final answer?

OpenStudy (anonymous):

That's what I'm trying to figure out and I can't seem to get it

OpenStudy (misty1212):

it is a straight forward substitution plug in the numbers as above that is all

OpenStudy (anonymous):

So y=(X-2)^2+1?

OpenStudy (misty1212):

no dear, the minus sign comes with the number

OpenStudy (anonymous):

So y=(x+2)^2+1?

OpenStudy (misty1212):

\[y=(x-(-2))^2+1\] but of course you don't want to write \(-(-2)\)

OpenStudy (misty1212):

yeah that one has vertex \((-2,1)\)

OpenStudy (anonymous):

Thank you! I actually understand it now!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]